# Thermal transport through a single trapped ion under strong laser illumination

### v1

@@ III Results
Having established sound connections between coherence and current, we now deepen our understanding of the current by studying how the trapped ion responds to varying the temperature gradient. It is usually expected that the magnitude of the current will be monotonic with the temperature bias, but this is not always the case. Some systems present what is called negative differential conductivity (NDC) yomo2005 ; elste2006 ; li2006 ; benenti2009 ; he2010 . In our system, we observe this phenomenon when we keep the temperature of the hot reservoir constant and we increase the temperature bias by lowering the temperature of the cold reservoir, as presented in Fig. 5. In this example, the motional reservoir was kept at temperature $T_{M}=5\nu$. Parameters used were $\delta=0.8\nu$ and $\Omega=1.5\nu$. Note the linear behaviour for temperatures larger than $T_{E}\approx 2\nu$, which is typical of the Fourier law joshi1993 ; lepri2003 ; manzano2012
$J\propto\Delta T.$ (14)
The linearity is broken for small temperatures of the reservoir coupled to the internal part, i.e., for stronger temperature bias. In particular, we observe nonmonotonicity for temperatures lower than $T_{E}\approx\nu$, which is the signature of NDC.
Figure 6: Ratio $Q$ as a function of $\delta$ and $\Omega$. For this plot, the motional and electronic reservoirs were set, respectively, at temperatures $T_{M}=5\nu$ and $T_{E}=0.5\nu$.
We can also investigate the appearance of this phenomenon for more general values of detuning $\delta$ and coupling constant $\Omega$. In order to do that, it is convenient to define a figure of merit for the NDC by means of the ratio
$Q:=\frac{\mathrm{max}\bigl{[}J_{SS}(T_{E})\bigr{]}}{J_{SS}(T_{E}=0)}.$ (15)
The graphic definitions of these quantities are depicted in Fig. 5. Basically, we have $Q=1$ when the current behaves monotonically with the temperature bias, and $Q>1$ otherwise. This is indeed the case displayed in Fig. 5. In Fig. 6 we present the values of the ratio $Q$ across the parameter space $\Omega\times\delta$. Again, here we set the motional reservoir at temperature $T_{M}=5\nu$, and calculated how $J_{SS}$ varies with $T_{E}$ for each pair $(\Omega,\delta)$, in order to compute $Q$. It can be seen that for small values of both $\delta$ and $\Omega$, we do not observe a significant amount of nonmonotonicity. For $\delta,\Omega\gtrsim 0.5\nu$ we observe an increase in the signatures of NDC. Again, the circular patterns tend to appear, but this time they are accompanied by oscillations in the value of $Q$. By fixing $\delta$, we observe oscillation in the value of $Q$ when approching and crossing the circular line. This is best noticed at the second circular sector.
The last aspect of transport we want to discuss in this setup is the onset of current rectification motz2018 ; simon2019 ; meher2020 ; kalantar2021 and its sensibility to variations in the system parameters. When the reservoirs are swapped, an alteration in the current direction is anticipated. Nevertheless, there might be an asymmetry in the current magnitudes, i.e., the current might flow more intensely in one direction. This is called current rectification. A quantifier of this asymmetry is found in the so-called rectification factor meher2020
@@
which is bounded by $\lvert R\rvert\leq 1$. When the currents flow symmetrically, we have $R=0$, while $\lvert R\rvert=1$ means complete rectification, i.e., the current flows only in one direction, like in a perfect diode.
@@
In Fig. 7, we present how the rectification factor depends on $\delta$ and $\Omega$. The temperature of the cold and hot reservoirs were set to $T_{C}=0.5\nu$ and $T_{H}=5\nu$, respectively. We define the current $J_{\rightarrow}$ as being the one when the cold reservoir is coupled to the electronic part, while $J_{\leftarrow}$ means the opposite, the electronic part being coupled to the hot reservoir. Once again, we observe the circular patterns show up. Despite being considerably asymmetric throughout the parameter ranges, we note that, close to the first circular sector, we have a clear bias towards negative rectification factors. In the second sector this also seem to be true, but with sharper changes in $R$ close to the circle line. In the third circle, however, the step size used in the numeric discretization did not give enough resolution for us to make a precise observation. In fact, as it can also be observed in the other figures, it appears that the variations in the features studied here become increasingly sharper for the outer circular sectors. This can ultimately constrain numerical studies for large integers $m$ in Eq. (12), since finer and finer discretization steps will be needed.
Overall, what we can observe in this last plot and the previous ones is that the system composed of a trapped ion interacting with a laser exhibits a remarkably rich behavior when subjected to a temperature gradient. The circular patterns and the entire phenomenology resulting from different coupling and detuning regimes make this system highly attractive for the investigation of the fundamentals of quantum mechanics and non-equilibrium thermodynamics neqt .
@@ IV Final remarks
We investigated quantum heat transport through a trapped ion in several regimes of the coupling strength $\Omega$ and ion-laser detuning $\delta$. The heat current was driven by a reservoir coupled to the electronic part of the ion, and another coupled to the motional part. Since we considered couplings well into the strong regime, the use of the dressed master equation (DME) formalism became necessary. The current, then, was calculated for the steady state of the DME.
We found that the heat current forms an intriguing circular pattern in the $\delta\times\Omega$ space. The current maxima fall close to the circles $\delta^{2}+\Omega^{2}=(m\nu)^{2}$, for an integer $m$ multiple of the trap frequency $\nu$. Interestingly, it was recently shown tassis2023 that in the first circular sector, $\delta^{2}+\Omega^{2}=\nu^{2}$, it is possible to approximate, in a unitarily rotated frame, the trapped ion Hamiltonian to that of a Jaynes-Cummings model. This could assist future investigations of this system, since the JCM is an exactly solvable model. To the best of our knowledge, no such approximations are yet known for the other circular sectors.
Furthermore, we also observed that the residual coherence in the steady state, calculated with the canonical basis, also follows a similar circular pattern. When the hot reservoir is coupled to the motional part, we observed that current maxima correlate to sudden drops in the residual coherence. While, for the case of the hot reservoir coupled to the electronic reservoir, we had sudden peaks of leftover coherence when the current was maximal.
Other properties we investigated were differential heat conductivity and heat rectification. For the first, it was observed that, in the vast majority of parameter ranges considered, the system presents negative differential heat conductivity when one of the reservoirs was brought close to zero temperature. This means that, in those regions, the current presents nonmonotonic behaviour, decreasing as the temperature gradient increases. As for the heat rectification, a large asymmetry in the current was observed throughout the parameter space when the reservoirs were swapped. This may inspire applications such as thermal rectifiers and thermal diodes within this pivotal quantum technology setup. Finally, our study lays the foundation for further investigations into heat transport across controlled quantum systems.
@@ Acknowledgements
T. T. acknowledges financial support from Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES, Finance Code 001). F. B. and F. L. S. acknowledge partial support from Brazilian National Institute of Science and Technology of Quantum Information (CNPq INCT-IQ 465469/2014-0). F. L. S. acknowledges partial support from Fundação de Amparo a Pesquisa do Estado de São Paulo (FAPESP) Process No. 2021/14135-1, CNPq (Grant No. 305723/2020-0), and CAPES/PrInt (Grant No. 88881.310346/2018-01).

### v2

@@ III Results
Having established sound connections between coherence and current, we now deepen our understanding of the current by studying how the trapped ion behaves in response to varying temperature gradients. While it is generally expected that the current magnitude will increase monotonically with temperature bias, this is not always the case. Some systems exhibit what is known as negative differential conductivity (NDC) yomo2005 ; elste2006 ; li2006 ; benenti2009 ; he2010 .
In the following, we set the temperature of the motional reservoir to $T_{M}=5\nu$ and vary the temperature of the reservoir coupled to the electronic part, $T_{E}$. In Fig. 7, we present results for a fixed detuning $\delta=0.8\nu$ and different values of the Rabi frequency $\Omega$. In all cases, we observe the onset of NDC, where the current no longer increases as the temperature gradient, $\Delta T\equiv T_{M}-T_{E}$, increases. However, for smaller temperature differences, the current exhibits the linear behavior $J\propto\Delta T$, consistent with Fourier’s Law. Similar NDC behavior is observed for different values of $\delta$ and $\Omega$, indicating a consistent response across these parameters. Additionally, when $\Delta T\approx 0$, the current behaves as expected, with heat flowing from the hot to the cold reservoirs, as indicated by the change in the signal of $J_{SS}$.
Figure 7: (Color online) Heat current $J_{SS}$ as a function of the temperature bias $\Delta T\equiv T_{M}-T_{E}$. For each line, the coupling strength was set to: $\Omega=0.4\nu$ (dashed grey); $\Omega=0.8\nu$ (solid purple); $\Omega=1.5\nu$ (dotted black); and $\Omega=1.9\nu$ (dash-dotted cyan) The detuning was set to $\delta=0.8\nu$. The temperature of the motional reservoir was fixed at $T_{M}=5\nu$.
Negative differential conductivity (NDC) is typically associated with the onset of nonlinearities. This well-established fact is exemplified by the well-known result that coupled harmonic oscillators cannot exhibit NDC li2006 ; he2010 . The results shown in Fig.(7) can thus be understood as arising from the inherent nonlinearity introduced by the laser-driven two-level system, which leads to the exponential function $\exp[\pm i\eta(a+a^{\dagger})]$ in the interaction part of the Hamiltonian, as given in Eq.(1). At sufficiently large temperature gradients, higher energy levels become populated, making the contribution of nonlinearity appreciable, leading to the observed NDC.
The next and final aspect of transport we examine in this work is the onset of current rectification motz2018 ; simon2019 ; meher2020 ; kalantar2021 and its sensitivity to variations in the same parameters as before, which are tuned through the control of the laser — specifically, the detuning $\delta$ and the coupling constant $\Omega$. When the reservoirs are swapped, the current direction reverses, but an asymmetry in current magnitude may emerge, with a stronger flow in one direction. This can be noted by comparing the magnitudes of $J_{SS}$ in Figs. 3a and 3b. A quantifier of this asymmetry is found in the so-called rectification factor meher2020
@@
where $J_{\rightarrow}$ is the current when the cold reservoir is coupled to the electronic part, and $J_{\leftarrow}$ is the current in the opposite scenario. From its definition in Eq. (14), we see that the rectification factor is bounded by $\lvert R\rvert\leq 1$. Specifically, $R=0$ indicates symmetric current flow, while $\lvert R\rvert=1$ represents complete rectification, where the current flows only in one direction, as in a perfect diode.
Up to this point, we have considered equal relaxation rates for both reservoirs, $\gamma_{Ejk}=\gamma_{Mjk}=\gamma$, in Eq.(6). Next, we examine the effects on the rectification factor of introducing different such rates. In Fig.8, we present the rectification factor $R$ as a function of the coupling strength $\Omega$, for a fixed detuning $\delta=0.8\nu$. The three curves shown correspond to a fixed electronic reservoir relaxation rate $\gamma_{Ejk}=\gamma$, while varying the relaxation rates for the motional reservoir. Specifically, we considered the cases $\gamma_{Mjk}\equiv\gamma_{M}=0.5\gamma$ (dashed grey), $\gamma_{M}=\gamma$ (solid purple), and $\gamma_{M}=2\gamma$ (dash-dotted cyan). The locations of the circular sectors are marked by dotted vertical lines. It can be seen that the rectification factor is more affected by changes in the relaxation rates in some regions than in others. In particular, in the first circular sector (first dotted vertical line), an increase in the relaxation rate $\gamma_{M}$ leads to more negative values of $R$. Conversely, in the third circular sector, a decrease in $\gamma_{M}$ facilitates negative rectification. There are also small regions where the curves cross, which remain robust to small changes in the relaxation rates.
This phenomenology suggests that the response of $R$ to variations in the thermal relaxation rates, combined with the flexibility offered by lasers of different intensities and detunings, opens up avenues for future investigations. These studies could focus on optimizing these parameters to achieve targeted values of the rectification factor. This approach could further broaden the potential applications of this system for thermal devices kalantar2021 and non-equilibrium thermodynamics neqt .
Figure 8: (Color online) Rectification factor as a function of the coupling constant $\Omega$ for different decay rates. For each line, the decay rate of the reservoir coupled to the motional part was set to: $\gamma_{M}=0.5\gamma$ (dashed grey); $\gamma_{M}=\gamma$ (solid purple); and $\gamma_{M}=2\gamma$ (dash-dotted cyan). The decay rate for the electronic reservoir was always set to $\gamma_{E}=\gamma$ The hot and cold reservoirs were fixed, respectively, at temperatures $T_{H}=5\nu$ and $T_{C}=0.5\nu$.
To gain a deeper understanding of how rectification depends on the parameters controlled by the laser properties, we once again fix the dissipation rates as $\gamma_{M}=\gamma_{E}=\gamma$ and present a simulation of the rectification factor for various values of $\delta$ and $\Omega$, as shown in Fig. 9. Again, circular patterns tend to emerge across the parameter space, but now the phenomenology is more complex. We observe a tendency toward negative rectification factors near the first circular sector. This trend appears in the second sector as well, though with sharper variations in $R$ close to the circular boundary. When approaching the third circular sector, however, the step size used in the numeric discretization did not have enough resolution for us to make a precise observation. In fact, as it can also be observed in the other figures, it appears that the variations in the features studied here become increasingly sharper for the outer circular sectors. This can ultimately constrain numerical studies for large integers $m$ in Eq. (12), since finer and finer discretization steps will be needed.
@@ IV Final remarks
We investigated quantum heat transport through a trapped ion in various regimes of the coupling strength $\Omega$ and ion-laser detuning $\delta$. The heat current was driven by one reservoir coupled to the electronic part of the ion and another coupled to the motional part. Given that we considered couplings well into the strong regime, the use of the dressed master equation (DME) formalism was necessary. The current was calculated for the steady state of the DME.
We found that the heat current forms an intriguing circular pattern in the $(\delta,\Omega)$ space. The current maxima fall close to the circles $\delta^{2}+\Omega^{2}=(m\nu)^{2}$, for an integer $m$, where $\nu$ is the trap frequency. Furthermore, we observed that the coherence in the steady state, calculated with the decoupled basis, exhibits similar circular patterns, but with a notable distinction: it is not always the maximum that aligns with the circular lines. When the hot reservoir is coupled to the motional part, we observed that current maxima correlate with sudden drops in the residual coherence. Conversely, when the hot reservoir is coupled to the electronic part, we observed sudden peaks of leftover coherence when the current was maximal.
Other properties we investigated included differential heat conductivity and heat rectification, where we observed a rich phenomenology as the coupling strength $\Omega$ was varied. This may inspire applications such as thermal rectifiers and thermal diodes within this pivotal quantum technology setup. Finally, our study lays the foundation for further investigations into heat transport across controlled quantum systems.
@@ Acknowledgements
T. T. acknowledges financial support from Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES, Finance Code 001). F. B. and F. L. S. acknowledge partial support from Brazilian National Institute of Science and Technology of Quantum Information (CNPq INCT-IQ 465469/2014-0). F. L. S. acknowledges partial support from Fundação de Amparo a Pesquisa do Estado de São Paulo (FAPESP) Process No. 2021/14135-1, CNPq (Grant No. 305723/2020-0) and (Grant No. 313068/2023-2), and CAPES/PrInt (Grant No. 88881.310346/2018-01).